# Fuzzy Prediction Sets: Conformal Prediction with E-values

### v2

@@ 3.2 Prediction as a testing problem
A second insight is that a prediction set $C_{\alpha}$ is equivalent to a hypothesis test. This is a powerful observation, because the construction of hypothesis tests is one of the most deeply studied topics in statistics.
Recall from the introduction that we may view a prediction set $C_{\alpha}$ for $\mathcal{P}$ as a map $C_{\alpha}:\mathcal{Z}^{n+1}\to\{0,1\}$, which returns an inclusion (0) or exclusion (1) for each point $z^{n+1}\in\mathcal{Z}^{n+1}$. We choose to interpret this map $C_{\alpha}$ as a test, where 0 represents a non-rejection and 1 a rejection. Coupling this to Figure 3, the white region corresponds to the non-rejection region of this test and the grey region to its rejection region.
The prediction set $C_{\alpha}$ (and by extension $C_{\alpha}^{Z^{n}}$) is valid for the model $\mathcal{P}$ if and only if it is a valid test for the ‘hypothesis’ $\mathcal{P}$, since
@@
A prediction set $C_{\alpha}^{Z^{n}}$ is equivalent to the test $C_{\alpha}$,
$C_{\alpha}(z^{n+1}):=\mathbb{I}\{z^{n+1}\not\in C_{\alpha}^{z^{n}}\}.$
The prediction set $C_{\alpha}^{Z^{n}}$ is valid at level $\alpha$ for the model $\mathcal{P}$ if and only if this test $C_{\alpha}$ is valid at level $\alpha$ for the hypothesis $\mathcal{P}$.
@@
At first glance the connection to testing may seem superficial, but we argue it is not. Recall that in classical hypothesis testing we hypothesize a model $\mathcal{P}$ (often called “the hypothesis”) and subsequently observe data. As we observe the data, we must reject the hypothesized model if the data and model are incompatible. In the construction of a prediction sets the roles are reversed: we fix a model $\mathcal{P}$ as the ground truth and hypothesize data $Z^{n+1}=z^{n+1}$. If the two are incompatible, we must now reject the hypothesized data. Extending this argument: if a part $Z^{n}=z^{n}$ of $Z^{n+1}$ is also observed alongside the model $\mathcal{P}$, we must reject the hypothesized part: $Z_{n+1}=z_{n+1}$.
@@ 3.3 Optimal prediction sets and power
By Theorem 1, constructing a prediction set is equivalent to constructing a test. It then remains to select a test that yields a ‘good’ or even optimal prediction set in some appropriate sense. One may hope that an optimal prediction set corresponds to a test that is most powerful against some alternative. We show this is indeed the case, by deriving the alternative hypothesis under which the test should be most powerful in order for our prediction set to be ‘as small as possible’.
@@
In Theorem 2, we show that $C_{\alpha}$ minimizes the expected size of $C_{\alpha}^{Z^{n}}$ if and only if its test-representation is most powerful against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu$. In the result, we even show that the choice of $\mu$ may depend on $Z^{n}$.

### v3

@@ 2.3 Prediction as a testing problem
A second insight is that a prediction set $C_{\alpha}$ is equivalent to a single hypothesis test. This is a powerful observation, because the construction of hypothesis tests is one of the most deeply studied topics in statistics.
Recall from the introduction that we may view a prediction set $C_{\alpha}$ for $\mathcal{P}$ as a map $C_{\alpha}:\mathcal{Z}^{n+1}\to\{0,1\}$, which returns an inclusion (0) or exclusion (1) for each point $z^{n+1}\in\mathcal{Z}^{n+1}$. We choose to interpret this map $C_{\alpha}$ as a test, where 0 represents a non-rejection and 1 a rejection. Coupling this to Figure 3, the white region corresponds to the non-rejection region of this test and the grey region to its rejection region. To avoid using two different names for the same object, we use the notation $C_{\alpha}$ both for the set and test representation.
The prediction set $C_{\alpha}$ for $Z^{n+1}$ (and by extension the prediction set $C_{\alpha}^{Z^{n}}$ for $Z_{n+1}$) is valid for the model $\mathcal{P}$ if and only if it is a valid test for the ‘hypothesis’ $\mathcal{P}$, since
@@
A prediction set $C_{\alpha}^{Z^{n}}$ is equivalent to the test $C_{\alpha}(z^{n+1}):=\mathbb{I}\{z^{n+1}\not\in C_{\alpha}^{z^{n}}\}$. The prediction set $C_{\alpha}^{Z^{n}}$ is valid at level $\alpha$ for the model $\mathcal{P}$ if and only if this test $C_{\alpha}$ is valid at level $\alpha$ for the hypothesis $\mathcal{P}$.
@@
At first glance the connection to testing may seem superficial, but we argue it is not. Recall that in classical hypothesis testing we hypothesize a model $\mathcal{P}$ (often called “the hypothesis”) and subsequently observe data. As we observe the data, we must reject the hypothesized model if the data and model are incompatible. In the construction of a prediction sets the roles are reversed: we fix a model $\mathcal{P}$ as the ground truth and hypothesize data $Z^{n+1}=z^{n+1}$. If the two are incompatible, we are forced to reject the hypothesized data. Extending this argument: if a part $Z^{n}=z^{n}$ of $Z^{n+1}$ is also observed alongside the model $\mathcal{P}$, we must reject the hypothesized part: $Z_{n+1}=z_{n+1}$.
@@ 2.4 Optimal prediction sets and power
By Theorem 1, constructing a prediction set is equivalent to constructing a test. It remains to select a test that yields a ‘good’ or even optimal prediction set in some appropriate sense. One may hope that an optimal prediction set corresponds to a test that is most powerful against some alternative. We show this is the case, by deriving the alternative hypothesis under which the test should be most powerful for a prediction set to be ‘as small as possible’.
@@
In Theorem 2, we show that $C_{\alpha}$ minimizes the expected size of $C_{\alpha}^{Z^{n}}$ if and only if its test-representation is most powerful against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu$. In the result, we even show that the choice of $\mu$ may depend on $Z^{n}$. The proof of this result and all other omitted proofs can be found in Appendix A. In Section 2.6, we include four examples where we apply this result in a Gaussian location setting.